# Fine-Tuning Flow Matching via Maximum Likelihood Estimation of Reconstructions

### v1

@@ Abstract
Flow Matching (FM) algorithm achieves remarkable results in generative tasks especially in robotic manipulation. Building upon the foundations of diffusion models, the simulation-free paradigm of FM enables simple and efficient training, but inherently introduces a train-inference gap. Specifically, we cannot assess the model’s output during the training phase. In contrast, other generative models including Variational Autoencoder (VAE), Normalizing Flow and Generative Adversarial Networks (GANs) directly optimize on the reconstruction loss. Such a gap is particularly evident in scenarios that demand high precision, such as robotic manipulation. Moreover, we show that FM’s over-pursuit of straight predefined paths may introduce some serious problems such as stiffness into the system. These motivate us to fine-tune FM via Maximum Likelihood Estimation of reconstructions - an approach made feasible by FM’s underlying smooth ODE formulation, in contrast to the stochastic differential equations (SDEs) used in diffusion models. This paper first theoretically analyzes the relation between training loss and inference error in FM. Then we propose a method of fine-tuning FM via Maximum Likelihood Estimation of reconstructions, which includes both straightforward fine-tuning and residual-based fine-tuning approaches. Furthermore, through specifically designed architectures, the residual-based fine-tuning can incorporate the contraction property into the model, which is crucial for the model’s robustness and interpretability. Experimental results in image generation and robotic manipulation verify that our method reliably improves the inference performance of FM.
@@ 1 Introduction
Deep generative models refer to a category of deep learning techniques designed to approximate and generate samples from an unknown underlying data distribution. A mainstream paradigm is to learn a mapping between a fixed (e.g., standard normal) distribution and the data distribution. This category notably includes diffusion models, which are the current state of the art on many gernerative modelling tasks. Particularly, they have also achieved remarkable results in robot motion generation tasks (Chi et al., 2023). The mathematical principles behind diffusion can be described by SDEs (Song et al., 2021). Naturally, we can also establish the relationship between noise and samples through ODE trajectories to simplify the model and achieve faster training and inference times. This inspired the development of the Flow Matching (FM) algorithm (Lipman et al., 2023; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023). FM has garnered extensive attention, particularly emerging as the leading approach in robot policy due to its fast inference speed (Black et al., 2024; Zhang et al., 2025; Braun et al., 2024; Chisari et al., 2024; Zhang & Gienger, 2024).
@@
FM, which inherits the characteristics of diffusion, employs a simulation-free training approach. This means that during the training phase, we only train some intermediate variables, e.g., vector filed (Lipman et al., 2023), score (Song & Ermon, 2019), and noise (or the previous state) (Ho et al., 2020). We cannot directly observe and optimize the final output from these difference or differential terms. In contrast, other generative models directly includes generated samples in their training loss. In Variational Autoencoder (VAE), we contain the rescontruction error (Kingma & Welling, 2013). In normalizing flow, we use the Maximum Likelihood Estimation (MLE) of the final output generated by the model (Rezende & Mohamed, 2015; Chen et al., 2018). Generative Adversarial Networks (GANs) are similar except they use adversarial training to replace the likelihood function (Goodfellow et al., 2014).
@@
Therefore, though offering relatively fast training speed, we have no knowledge of how the real samples are actually generated during the simulation-free training phase. This implies the existence of a gap between its training and inference phases. Such a gap can considerably impact scenarios that require high precision, such as robotic manipulation or spatio-temporal data imputation and prediction. By comparison, the Action Chunking with Transformers (ACT) architecture (Zhao et al., 2023), built upon a VAE with integrated reconstruction error, performs remarkably well in fine manipulation tasks. Moreover, we point out that FM’s over-pursuit of straight paths may render the system stiff or even lead to discontinuous vector fields (see Fig. 1(a)), resulting in numerical instability and model unreliability. Although in practice, the stochastic noise introduced by batching and early stopping technique can smooth the model output and mitigate this phenomenon, it comes at the cost of underfitting. Fortunately, due to the smoothness of ODE trajectories (Chen et al., 2018), it is feasible to fine-tune FM directly by reconstruction error. There are multiple ways to track parameter gradients (Kidger, 2021, Chapter 5) such as adjoint sensitivity method.
@@
The schematic diagrams in Figure 1 outline the core principles of our proposed methodology. Fine-tuning enhances the representational capacity of vector fields while preserving their simplicity. Notably, it can imbue a flow with contraction properties, leading to robust stability against minor perturbations and a latent space with higher semantic quality. This paper is organized as follows. We first theoretically analysis the relation between training loss and inferring error (Theorem 1). Then we propose a method of fine-tuning FM via MLE of reconstruction inference error (Theorem 2), which includes both straightforward fine-tuning and residual-based fine-tuning approaches. Furthermore, through specifically designed architectures, the residual-based fine-tuning can incorporate contraction properties into the model (Theorems 3-4). Experimental results verify that our method reliably improves the inference performance of FM. Our primary contributions are: (1) the first theoretical analysis quantifying the relationship between training and inference errors in FM; (2) a practical MLE-based fine-tuning framework that includes both an easy-to-implement version and a robustness-oriented residual variant with contraction analysis; and (3) comprehensive experimental validation.
@@
Figure 1: Fine-Tuning Your Flow: A Visual Explanation. These figures plot trajectories under different vector field models. (a) illustrates that over-pursuing straightness (blue lines) leads to discontinuities in the vector field, i.e., $f(0^{+},0)\neq f(0^{-},0)$. This will cause the system to exhibit stiffness, significantly exacerbating the difficulty of numerical solution and thereby compromising the model’s reliability. By comparison, the green line depicts a more stable flow. (b) plots the pre-trained FM path (blue line), the fine-tuned flow (green line), the flow fine-tuned with residuals (red line), and the flow trained entirely with MLE (purple line). An oversimplified assumption like "the straight path" can lead to underfitting (blue line). A fine-tuned model converges to a local optimum near the pre-trained model, thereby improving its fit to the sample while maintaining path simplicity (green line). By contrast, the CNF (purple line) solely fits the samples without any prior guidance on the vector field’s shape, which can easily produce overly complex trajectories. The red line utilizes a residual fine-tuning approach that preserves the pre-trained model unchanged, employing solely a residual network to learn the remaining residual components. (c) plots the variation curve of component $x_{1}$ over time. Here, compared with the blue line, the green one represents a "contracting" trajectory (Lohmiller & Slotine, 1998). When subjected to a minor disturbance $d_{0}$ (which may arise from stochastic noise or slight differences in external inputs), a contracting trajectory still tends to stabilize around a similar solution. Such a model demonstrates superior performance in terms of stability and robustness. (d) illustrates contraction in 2-dimensional space. (Blue) points in different contraction regions will converge to different (red) destinations. Points in the same contraction region behave stably and robustly.
@@ 2 Related work
Neural Ordinary Differential Equations (NODE), or Continuous Normalizing Flow (CNF) in the context of generation tasks (Chen et al., 2018), also adpot the ODE model. In contrast to FM’s dependence on predefined reference paths, these methods utilize an optimization framework grounded in Maximum Likelihood Estimation (MLE), using the change of variables theorem with a tractable noise distribution to compute probabilities. Consequently, they tend to generate complex and intractable vector fields (see Fig. 1(b)). Significant efforts have been devoted to its improvement. Dupont et al. (2019) project the state into a higher-dimensional space to enable simpler paths. Finlay et al. (2020) introduce regularizations that encourage neural ODEs to prefer simpler dynamics. However, in most tasks, these methods still fail to match the simplicity and efficiency of simulation-free approaches like diffusion and FM. But their direct optimization of the end-to-end loss theoretically allows for a higher potential performance.
Convergence and Stability
Some work is also being done on the progressive stability analysis of networks based on ODE models. Llorente-Vidrio et al. (2021) study a class of NODEs, using the asymptotic stability to guide the design of network weights. Mei et al. (2024) investigate the theoretical conditions for convergence. These two studies are primarily focused on the task of image classification by NODE, though Mei et al. (2024) experimentally explore the ability of their network to model dynamics. The theoretical analysis of convergence, stability and contraction focus on the final steady-state characteristics of the model, rather than the preliminary dynamic behavior. Our work is for the flow model in generation tasks (one of the most prevalent approaches in this domain especially in robot manipulation). Generative models desires solution diversity, but convergence (Mei et al., 2024; 2022; Efimov & Aleksandrov, 2021) refers to a global convergent (unique) solution. Thus, in this paper we study contraction (Lohmiller & Slotine, 1998) instead, which is more concerned with the local aspect: do nearby trajectories converge to one another? This paper enables the partitioning of the domain into distinct contraction regions and then analyzes the contraction within each one. Moreover, in terms of network design, we adhere to the principle that dynamic characteristics (as long as the state remains bounded) do not affect the subsequent steady-state contraction property. We simply added a residual fine-tuning network segment at the later stage. This approach not only preserves the powerful approximation capabilities of the flow-matching component involving UNet or Transformer structures but also ensures stability and contraction in the later phases. Note that stability analyses have also been performed on NODE variants including SODEF (Kang et al., 2021), SNDEs (White et al., 2023), and Stable Neural Flows (Massaroli et al., 2020).
Optimal-transport conditional FM
@@
The original FM method constructs paths based on individual sample pairs, which may lead to twisted, entangled, and complex trajectories across the overall distribution. Several studies have been dedicated to addressing this issue. Liu (2022) propose Rectified Flow, which leverages the observation that flow models follow a global predefined path yet make local adjustments to noise-sample pairings. This enables trajectory distillation but suffers from accumulated error through iteration. Tong et al. (2024) and Pooladian et al. (2023) consider the joint probability between noise and samples within a batch. By constructing optimal transport within minibatches, it approximates the optimal path between global distributions. This paper applies the principles of Tong et al. (2023; 2024) to determine the pairing between noise and samples.
Consistency and Fast Inference
The consistency model (Song et al., 2023) or consistency FM (Yang et al., 2024) represents another research direction aimed at achieving rapid single-step inference, albeit with a deliberate compromise on model accuracy. The Consistency Model learns to map any point on a PF-ODE trajectory directly to its origin (clean data). The reliance on distillation limits their practicality in robotics (Lu et al., 2024). Consistency flow defines a straight flow from any time step to a fixed endpoint by constraining its velocity, and has been applied in robot policy (Zhang et al., 2025). However, as previously discussed, these practices can lead to an excessively large Lipschitz constant (Fig. 1(a)), resulting in stiff behavior. This could pose potential risks for a system described by differential equations. Our vision is to accelerate inference speed by constructing a simple, smooth, and contracting vector field that fundamentally reduces the difficulty of numerical solutions. Furthermore, with the help of contraction analysis, we could possibly map each contraction region to its corresponding equilibrium point, thereby achieving single-step inference without relying on distillation and without loss of accuracy.
3 On the Fine-Tuning of FM: Motivations and Methods
3.1 The Training-Inference Gap in FM
@@
Given two marginal distributions $q_{0}(x_{0})$ and $q_{1}(x_{1})$ for which we would like to learn a model to transport between, FM seeks to optimize the simple regression objective $\mathbb{E}_{t,p_{t}(x)}\left\|v_{t}(x;\theta)-u_{t}(x)\right\|^{2}$, where $u_{t}(x)$ is the is a vector field that generates a probability path $p_{t}$ under the two marginal constraints, $v_{t}(x;\theta)$ is the parametric vector field. Let $\phi_{t}(x_{0})$ be the solution to the ODE $\frac{d}{dt}\phi_{t}=v_{t}(\phi_{t};\theta)$ with initial value $\phi_{0}=x_{0}$. To obtain the numerical solution, let us define $N$ as the total number of discrete steps, $t_{i}$ as the $i$-th time point, $\tau_{i}=t_{i+1}-t_{i}$ as $i$-th interval. Different time points typically involve different step sizes if we employ adaptive step size algorithms, as is often the case. We use $\hat{\phi}_{n}(x_{0})$ to represent the corresponding numerical solution at time $t_{n}$, with the model error $\varepsilon_{n}(x_{0}):=\psi_{t_{n}}(x_{0})-\hat{\phi}_{n}(x_{0})$.
@@
For computational tractability, we use the equivalent Conditional Flow Matching (CFM) objective $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{t}(x|x_{1})}\left\|v_{t}(x)-u_{t}(x|x_{1})\right\|^{2}$ (Lipman et al., 2023). Reparameterizing $p_{t}(x|x_{1})$ in terms of just $x_{0}$ we get
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t,q(x_{0},x_{1})}\left\|v_{t}\left(\psi_{t}(x_{0}|x_{1});\theta\right)-u_{t}\left(\psi_{t}(x_{0}|x_{1})|x_{1}\right)\right\|^{2},$ (2)
where $\psi_{t}(x_{0}|x_{1})$ is the conditional flow with a predefined form. We typically use $\psi_{t}(x_{0}|x_{1})=(1-t)x_{0}+tx_{1}$ with the corresponding vector field $u_{t}\left(\psi_{t}(x_{0}|x_{1})|x_{1}\right)=x_{1}-x_{0}$. Thus we have
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t,q(x_{0},x_{1})}\left[\|x_{0}+\Delta tv_{t}(t,\psi_{t}(x_{0}|x_{1});\theta)-x_{1}\|^{2}\right],\quad\text{with}\ \Delta t=1.$ (3)
@@
This can be seen as measuring the ground truth $x_{1}$ and the numerical result by implicit one-step Euler method within time interval $[0,1]$, with derivative estimated in time $t$. We can also turn to a more advanced solving scheme, as we do during the inference phase. The difference is that when in training we use the ground truth value $\psi_{t}(x_{0}|x_{1})$ since we have predefined the path, but when inferring we should use the estimated value $\hat{\phi}_{n}(x_{0})$. And this creates the gap between the training stage and inference (or prediction). Fortunately, we can bound this gap by the following theorem.
Theorem 1.
@@
Assume that the truth vector field $u_{t}(x)$ is a Lipschitz-continuous function with the Lipschitz constant $L_{u}$ > 0. And the discrepancy between the learned vector field and the truth satisfies $\|v_{\theta}(t,x)-u_{t}(x)\|_{\infty}\leq\delta$, then we can derive the following error estimate between the ground-truth values and the network’s inferred values
$|\varepsilon_{N}|\leq\exp{(L_{u}t_{N-1})}\left(\Sigma_{j=0}^{N-1}(\delta\tau_{j}+\frac{1}{2}M\tau_{j}^{2})+|\varepsilon_{0}|\right).$ (4)
@@
where $M=\max\limits_{0\leq t\leq 1}|\ddot{\psi_{t}}|$ is an upper bound for the second time derivative. Under special circumstances, when uniform step sizes are adopted, we obtain a more refined estimation formula,
$|\varepsilon_{N}|\leq\exp({L_{u}})\varepsilon_{0}+\frac{M\tau_{0}+2\delta}{2L}(\exp{(L_{u}})-1).$ (5)
Proof is in Appendix B.
Remark 1.
Remark 2.
Theorem 1 generalizes classical numerical analysis to settings in which the underlying vector field is approximated (e.g., via learning), introducing inherent approximation errors.
@@
Though this shows that the gap between training and inference is bounded, such a gap will inevitably compromise the model’s effectiveness. This theorem indicates that the error in the final generated sample (or action in robot policy) $\varepsilon_{N}$ will further amplify the training error $\delta$, at least by a multiplicative factor $\exp({L_{u}})$ with an additive constant. This issue becomes particularly severe when predefined paths may cause discontinuities in the vector field (Fig. 1(a)), resulting in a large Lipschitz constant $L_{u}$. This motivates our pursuit of a consistent training-inference paradigm that can further optimaize $\varepsilon_{N}$. More specifically, we utilize MLE for fine-tuning based on reconstruction results, and the detailed principles and procedures will be thoroughly discussed in the following subsection.
3.2 MLE Fine-tuning

### v2

@@ Abstract
Flow Matching (FM) models achieve remarkable results in generative tasks. Building upon diffusion models, FM’s simulation-free training paradigm enables simplicity and efficiency but introduces a train-inference gap: model outputs cannot be assessed during training. Moreover, the straight flow assumption suffers from some inherent limitations. To address this, we propose to fine-tune FM via Maximum Likelihood Estimation (MLE) of reconstructions—enabled by FM’s smooth ODE formulation, unlike the stochastic differential equations (SDEs) in diffusion models. We first theoretically analyze the relationship between training loss and inference error in FM under numerical precision constraints. We then propose an easy-to-implement fine-tuning framework based on MLE of reconstructions, with flexibility for sophisticated extensions. Building on this, we incorporate a generalized artificial viscosity term that enhances flow stability and robustness, accompanied by a direct parameterization method and rigorous theoretical guarantees. Experiments demonstrate our method’s effectiveness across diverse settings: a toy example provides mechanistic insights into the fine-tuning process, while large-scale evaluations on meteorological forecasting and robotic manipulation policies validate reliable performance improvements.
Machine Learning, ICML
@@ 1 Introduction
Deep generative models refer to a category of deep learning techniques designed to approximate and generate samples from an unknown underlying data distribution. A mainstream paradigm is to learn a mapping between a fixed (e.g., standard normal) distribution and the data distribution. This category notably includes diffusion models, a leading approach for many generative tasks across vision (Rombach et al., 2022; Ho et al., 2022), policy (Chi et al., 2023), and spatio-temporal data (Li et al., 2022; Yuan et al., 2023). The mathematical principles behind diffusion can be described by SDEs (Song et al., 2021b). Naturally, we can also establish the relationship between noise and samples through ODE trajectories to simplify the model. This inspired the development of the Flow Matching (FM) algorithm (Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023). FM has garnered extensive attention (Irvin et al., 2025; Rohbeck et al., 2025; Wang et al., 2025; Esser et al., 2024; Hamad and Rosenbaum, 2025), particularly emerging as a mainstream approach in robot policy due to its fast inference speed (Black et al., 2024; Zhang et al., 2025; Braun et al., 2024; Chisari et al., 2024; Zhang and Gienger, 2024).
@@
FM, which inherits the characteristics of diffusion, employs a simulation-free training approach. This means that during the training phase, we only train some intermediate variables, e.g., vector filed (Lipman et al., 2023), score (Song and Ermon, 2019), and noise (or the previous state) (Ho et al., 2020). We cannot directly observe and optimize the final output from these difference or differential terms. In contrast, other generative models directly includes generated samples in their training loss. In Variational Autoencoder (VAE), we contain the rescontruction error (Kingma and Welling, 2013). In normalizing flow, we use the Maximum Likelihood Estimation (MLE) of the final output generated by the model (Rezende and Mohamed, 2015; Chen et al., 2018). Generative Adversarial Networks (GANs) are similar except they use adversarial training to replace the likelihood function (Goodfellow et al., 2014).
@@
Therefore, though offering relatively fast training speed, we have no knowledge of how the real samples are actually generated during the simulation-free training phase. This implies the existence of a gap between its training and inference phases. Such a gap can considerably impact scenarios that require high precision, such as robotic manipulation or spatio-temporal numerical tasks. By comparison, the Action Chunking with Transformers (ACT) architecture (Zhao et al., 2023), built upon a VAE with integrated reconstruction error, performs remarkably well in fine manipulation tasks. Moreover, we point out that FM’s over-pursuit of straight paths may render the system stiff or even lead to discontinuous vector fields (see Figure 1(a)), resulting in numerical instability and model unreliability. Although in practice, the stochastic noise introduced by batching and early stopping technique can smooth the model output and mitigate this phenomenon, it comes at the cost of underfitting. Fortunately, due to the smoothness of ODE trajectories (Chen et al., 2018), it is feasible to fine-tune FM directly by reconstruction error. There are multiple ways to track parameter gradients (Kidger, 2021, Chapter 5) such as adjoint sensitivity method. In contrast, the SDE-based diffusion models can only estimate—not directly optimize—the maximum likelihood (Song et al., 2021a).
@@
The schematic diagrams in Figure 1 outline the core principles of our proposed methodology. Fine-tuning enhances the representational capacity of vector fields while preserving their simplicity. Notably, owing to the flexibility of the MLE fine-tuning framework, we can incorporate additional designs to induce models with favorable properties. Inspired by Scientific Computing, we introduce a generalized viscosity during fine-tuning, which endows the flow with enhanced stability and robustness against minor perturbations (i.e., contraction, a concept from control theory). This paper is organized as follows. In Section 3 we theoretically analysis the relation between training loss and inferring error . Then in Section 4, we propose a framework of fine-tuning FM via MLE of reconstruction error, and introduce an generalized viscosity for robustness. Section 5 shows the experimental results.
Our primary contributions are: 1) the first theoretical analysis of training-inference error in FM under numerical accuracy constraints; 2) an easy-to-implement fine-tuning framework centered on a concrete MLE loss, and a generalized artificial viscosity term within this framework to enhance robustness; 4) rigorous theoretical proofs for all proposed methods; and 5) comprehensive experiments from toy examples to large-scale meteorological and robotics tasks.
@@
Figure 1: Fine-Tuning Your Flow: A Visual Explanation. (a) illustrates that over-pursuing straightness (blue lines) leads to discontinuities in the vector field, i.e., $f(0^{+},0)\neq f(0^{-},0)$. By comparison, the green line depicts a more stable flow. (b) plots the pre-trained FM path (blue line), the fine-tuned flow (green line), the flow fine-tuned with residuals (red line), and the flow trained entirely with MLE (purple line). An oversimplified assumption like ”the straight path” can lead to underfitting (blue line). A fine-tuned model converges to a local optimum near the pre-trained model, thereby improving its fit to the sample while maintaining path simplicity (green line). By contrast, the CNF (purple line) often fits the samples with overly complex trajectories. The red line utilizes a residual network to learn the remaining residual components. (c) plots the variation curve of component $x_{1}$ over time. Here, compared with the blue line, the green one represents a “contracting” trajectory. When subjected to a minor disturbance $d_{0}$ , a contracting trajectory still tends to stabilize around a similar solution, showing superior stable performance. (d) illustrates contraction in 2-dimensional space. (Blue) points in different contraction regions will converge to different (red) destinations. Points in the same contraction region behave stably and robustly.
@@ 2 Related work
Neural Ordinary Differential Equations (NODE), or Continuous Normalizing Flow (CNF) in the context of generation tasks (Chen et al., 2018), also adopt the ODE model. In contrast to FM’s dependence on predefined reference paths, these methods utilize an optimization framework grounded in MLE, using the change of variables theorem with a tractable noise distribution to compute probabilities. Consequently, they tend to generate complex and intractable vector fields. Significant efforts have been devoted to its improvement (Dupont et al., 2019; Finlay et al., 2020). However, these methods still fail to match the efficiency of simulation-free approaches like FM.
Optimal-Transport Conditional FM.
@@
The original FM method constructs paths based on individual sample pairs, which may lead to twisted, entangled, and complex trajectories across the overall distribution. Several studies (Liu, 2022; Tong et al., 2024; Pooladian et al., 2023) have been dedicated to addressing this issue. The latter two consider the joint probability between noise and samples within a batch and approximate the optimal path between global distributions. Given the complexity of MLE training, this step is important for enhancing the performance of our method and accelerating its convergence.
Stability of ODE Model.
Current investigations regarding the stability of Neural ODEs are largely confined to the domain of classification (Llorente-Vidrio et al., 2021; Mei et al., 2024; Kang et al., 2021; Massaroli et al., 2020), while Mei et al. (2024) experimentally explore the ability of their network to model dynamics. Some works also address stability analysis in dynamic modeling, such as scenarios involving prior physical constraints (White et al., 2023) or external inputs (Li et al., 2025). Our work primarily focuses on the generative domain, encompassing prediction or decision-making tasks, which also represent the mainstream approaches for these tasks nowadays.
Fast Inference of FM.
Another popular research direction aims at improving inference speeds, achieving comparable inference quality in single or few steps to that of multi-step inference. Some methods rely on distillation (Geng et al., 2023; Salimans and Ho, 2022; Sauer et al., 2024; Luo et al., 2023; Yin et al., 2024; Zhou et al., 2024) which inevitably incurs the cost of distillation error. Other works impose constraints directly on the vector field (Song et al., 2023; Song and Dhariwal, 2023; Geng et al., 2025; Lu and Song, 2025; Yang et al., 2024). However, from the perspective of dynamical systems, such methods tend to increase system stiffness or even discontinuities. Our work amis to enhance the upper bound of the model’s inference capability by constructing a better-posed vector field.
3 Motivations of Fine-Tuning FM
3.1 Preliminaries and Notation
@@
Here $\bm{o}$ is the outside information or condition, and we may omit it when there is no ambiguity. Given two marginal distributions $q_{0}(x_{0})$ and $q_{1}(x_{1})$ for which we would like to learn a model to transport between, FM seeks to optimize the simple regression objective $\mathbb{E}_{t,p_{t}(x)}\left\|v_{t}(x;\theta)-u_{t}(x)\right\|^{2}$, where $u_{t}(x)$ is the is a vector field that generates a probability path $p_{t}$ under the two marginal constraints, $v_{t}(x;\theta)$ is the parametric vector field. Let $\phi_{t}(x_{0})$ be the solution to the ODE $\frac{d}{dt}\phi_{t}=v_{t}(\phi_{t};\theta)$ with initial value $\phi_{0}=x_{0}$. To obtain the numerical solution, let us define $N$ as the total number of discrete steps, $t_{i}$ as the $i$-th time point, $\tau_{i}=t_{i+1}-t_{i}$ as $i$-th interval. Different time points typically involve different step sizes if we employ adaptive step size algorithms, as is often the case. We use $\hat{\phi}_{n}(x_{0})$ to represent the corresponding numerical solution at time $t_{n}$, with the model error $\varepsilon_{n}(x_{0}):=\psi_{t_{n}}(x_{0})-\hat{\phi}_{n}(x_{0})$.
@@
For computational tractability, we need the equivalent Conditional Flow Matching (CFM) objective $\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{t}(x|x_{1})}\left\|v_{t}(x)-u_{t}(x|x_{1})\right\|^{2}$ (Lipman et al., 2023). Reparameterizing $p_{t}(x|x_{1})$ in terms of just $x_{0}$ we get $\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t,q(x_{0},x_{1})}\big\|v_{t}\left(\psi_{t}(x_{0}|x_{1});\theta\right)\ \ -u_{t}\left(\psi_{t}(x_{0}|x_{1})|x_{1}\right)\big\|^{2},$ where $\psi_{t}(x_{0}|x_{1})$ is the conditional flow with a predefined form. We typically use $\psi_{t}(x_{0}|x_{1})=(1-t)x_{0}+tx_{1}$ with the corresponding vector field $u_{t}\left(\psi_{t}(x_{0}|x_{1})|x_{1}\right)=x_{1}-x_{0}$. Thus we obtain the loss
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{t,q(x_{0},x_{1})}\big\|x_{0}+v_{t}(t,$ $\psi_{t}(x_{0}|x_{1});\theta)$ (2)
$\qquad\quad-\ x_{1}\big\|^{2}.$
3.2 The Training-Inference Gap
@@
Loss (2) can be seen as measuring the ground truth $x_{1}$ and the numerical result by implicit one-step Euler method within time interval $[0,1]$, with derivative estimated in time $t$. We can also turn to a more advanced solving scheme, as we do during the inference phase. The difference is that when in training we use the ground truth value $\psi_{t}(x_{0}|x_{1})$ since we have predefined the path, but when inferring we should use the estimated value $\hat{\phi}_{n}(x_{0})$. And this creates the gap between the training stage and inference (or prediction). Fortunately, we can bound this gap by the following theorem (proof in Appendix A.).
Theorem 3.1 (Training-Inference Gap).
@@
Assume that the truth vector field $u_{t}(x)$ is a Lipschitz-continuous function with the Lipschitz constant $L_{u}>0$. And the discrepancy between the learned vector field and the truth satisfies $\mathbb{E}_{p(x_{0})}\|v_{\theta}(t,\psi_{t}(x_{0}))-u_{t}(\hat{\phi}_{t}(x_{0}))\|_{\infty}\leq\delta$, then we can derive the following error estimate between the ground-truth values and the network’s inferred values
$\mathbb{E}\,|\varepsilon_{N}|\leq\exp{(L_{u}t_{N-1})}\left(\Sigma_{j=0}^{N-1}(\delta\tau_{j}+\frac{1}{2}M\tau_{j}^{2})+\mathbb{E}\,|\varepsilon_{0}|\right),$
@@
where $M=\max\limits_{0\leq t\leq 1}|\ddot{\psi_{t}}|$ is an upper bound for the second time derivative. Under special circumstances, when uniform step sizes are adopted, we obtain a more refined estimation,
$\mathbb{E}\,|\varepsilon_{N}|\leq\exp({L_{u}})\,\mathbb{E}\,|\varepsilon_{0}|+\frac{M\tau_{0}+2\delta}{2L}(\exp{(L_{u}})-1).$
Remark 3.2 (Theoretical Significance and Connections).
This theorem serves as a diagnostic tool, offering an end-to-end decomposition of the inference error into components attributable to the training loss, the vector field network’s properties ($L_{u}$), and numerical discretization. By explicitly quantifying the training-inference gap along the entire pipeline, it directly motivates our subsequent work on optimizing based on inference error (3). Compared to Benton et al. (2024), our work further accounts for the accuracy of numerical methods—a factor that significantly affects the model’s practical inference performance, as demonstrated in Figures 2, 3, and C.1. Approximating the underlying vector field (e.g., through learning) introduces inherent errors. The resulting question of stability and error amplification for classical numerical methods is the core theoretical problem that our theorem resolves. While Zhou and Liu (2025) also conduct error analysis, it is more focused on statistical learning theory, revealing the relationship between convergence rate and sample size. It does provide some numerical error analysis, but this belongs to the category of pure numerical analysis, quantifying the isolated discretization error of the approximate network itself rather than offering an end-to-end analysis from training loss to inference error. Its primary purpose is to guide how to balance computational cost and accuracy during inference.
@@
Though this shows that the gap between training and inference is bounded, such a gap will inevitably compromise the model’s effectiveness. This theorem indicates that the error in the final generated sample (or action in robot policy) $\varepsilon_{N}$ will further amplify the training error $\delta$, at least by a multiplicative factor $\exp({L_{u}})$ with an additive constant. This issue becomes particularly severe when predefined paths may cause discontinuities in the vector field (Figure 1(a)), causing $L_{u}\to\infty$. This motivates our pursuit of a consistent training-inference paradigm that can further optimaize $\varepsilon_{N}$. More specifically, we will utilize MLE for fine-tuning based on reconstruction results in the next section.
Remark 3.3 (Discontinuity).
Discontinuity differs between FM (ODE) and diffusion models (SDE): SDE discontinuity stems from its (Brownian) diffusion term following statistical laws, which does not exhibit stiff behavior.
4 MLE Fine-Tuning Framework for FM
In this section, we will introduce the specific methods for fine-tuning. We first introduce the basic fine-tuning framework, and then we investigate a way to incorporate more designs into this framework to enhance its robustness.
4.1 The Basic Framework